Calculus I, Fall 2012 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 14

1. Section 7.4: Problems: 27, 46, 55, 57.

2. Section 7.8: Read page 499-500, examples 1-2. Then do problems: 3, 4, 10.

3. Section 7.8: Do example 2 by slicing in the z-direction.